# Bhopal 462 066, Madhya Pradesh, India

### v1

@@ Abstract
The mechanical response of intrinsically disordered proteins (IDPs) and polyampholyte (PA) chains is vital for understanding their biological functions and designing functional materials. We investigate the force–extension behavior of a PA chain with distinct charge sequences using molecular dynamics simulations and a theoretical approach based on the generalized random-phase approximation (GRPA). A diblock PA chain under extensional force undergoes a continuous coil-to-stretch transition at weak electrostatic coupling, which sharpens into a globule–coil-like transition at stronger coupling. The GRPA theory quantitatively captures these behaviors, including the sharp conformational transition and its dependence on electrostatic strength. Simulations reveal pronounced hysteresis during the force-extension and relaxation processes. Additionally, the elastic modulus exhibits four regimes: an initial plateau, stress stiffening, an exponential stress-softening behavior, and a stress stiffening regime. Using the theoretical model and structural input of the PA chain, we have demonstrated that the elastic modulus in the elastic softening regime decreases exponentially, $E\sim\exp(-\alpha_{0}f/\Gamma_{e})$, as a function of $f$, which aligns with the simulation results. The elastic response of the PA chain is further examined across different charge sequences, where both elastic softening and sharp transitions are absent at smaller block lengths. Finally, coarse-grained models of IDPs such as LAF-1 and DDX4 exhibit similar nonlinear elasticity, highlighting the universality of these mechanisms. Our results establish a fundamental link between electrostatic correlations, charge sequence, and nonlinear elasticity, bridging molecular interactions and macroscopic mechanics.
keywords:
American Chemical Society,
1 Introduction
@@
A wide variety of biopolymers—including nucleic acids, and intrinsically disordered proteins (IDPs)—as well as synthetic macromolecular systems, contain both acidic and basic functional groups 1, 2, 3. These systems play essential roles in numerous biological processes, such as cell signaling 4, 5, transcriptional regulation 6, 7, muscle elasticity 8, liquid–liquid phase separation (LLPS), and the formation of biomolecular condensates 9, 10, 11, 12, 13, 14. Understanding their mechanical response is crucial for elucidating the behavior of biological networks 15, protein functionality 4, 5, 6, 7, and the folding energy landscapes of RNA and DNA 16, 17, 18, 19. Many such polymers and proteins possess folded structural domains that undergo cooperative unfolding under applied forces, resulting in complex force–extension curves (FECs) characterized by abrupt extensions and nonlinear mechanical responses 20, 21, 22, 23, 24.
@@
The theoretical frameworks for the elastic response of a polymer have led to interesting scaling relationships that describe mechanical behavior across different regimes25, 15, 26, 27, 28, 25, 29, 18, 30. These relationships have been supported by both computational approaches and experimental observations18, 31, 32, 33, 34, 35, 36, 20, 21, 22, 23, 24. The single‑molecule techniques such as atomic force microscopy (AFM) and optical/magnetic tweezers have significantly transformed the study of the polymer’s elastic behavior37, 38, 17, 34, 16, 39, 40, 41, 42, 18, 34, 43. These methods enable direct measurement of force–extension relationships at the level of individual chains, thereby validating theoretical predictions, identifying distinct mechanical regimes, and highlighting how molecular heterogeneity influences bulk mechanical behavior.
@@
The elastic response of a linear, uncharged polymer in the low-force regime is characterized by purely entropic behavior, with the force–extension curve being linear, and the elastic modulus ($E$) remaining constant. At large stretching forces, a nonlinear relationship emerges between force and extension. In the case of the freely jointed chain (FJC) model44, the extension approaches $R_{e}>\approx R_{max}[1-k_{B}T/f\ell_{K}]$, where $R_{e}$, $R_{max}$, $k_{B}T$, and $\ell_{K}$ are the chains extension, contour length, thermal energy, and Kuhn length, respectively. Alternatively, for semiflexible chains such as double-stranded DNA, a worm-like chain (WLC) model developed by Marko and Siggia provides the force extension as $R_{e}\approx R_{max}[1-\sqrt{k_{B}T/2f\ell_{K}]}$ 25, 45, 38, 46, 19, 47, 48. The WLC model provides a fairly good description of the elastic response of ssRNA, ssDNA, PolyEthylene Glycol, and various other proteins across a broad range of forces34, 49, 50, 18, 51, 15. In these large-force limits, the elastic modulus scales as $E\sim f^{2}$ for the FJC and $E\sim f^{1.5}$ for the WLC model52. More importantly, in the intermediate regimes, the force extension typically depends on the nature of the interactions among various domains of the polymer, excluded volume, electrostatic interactions, and ionic concentrations. These interactions can lead to distinct elastic behaviors.
@@
Previous investigations have examined the force–extension behavior of neutral polymers, polyelectrolytes, and zwitterionic polymers 35, 53, 54, 32, 33, 55. These models are typically formulated for homogeneously charged chains and therefore do not capture the complex electrostatic environments that can give rise to the rugged energy landscapes often observed in proteins. Despite extensive work on neutral and homogeneously charged polymers, the role of electrostatic sequence heterogeneity and charge segregation in determining the nonlinear elastic response, hysteresis, and softening behavior of polyampholytes (PAs) remains poorly understood. PA serves as a simplified model system for intrinsically disordered proteins (IDPs) and other charged proteins. Their conformational and phase behaviors depend intricately on parameters such as net charge, charge sequence or block length, chain rigidity, and ionic concentration 56, 57, 58, 59, 60, 61, 62. The influence of charge sequence and the strength of electrostatic interactions on their elastic behavior remains crucial yet not explored. Addressing this gap could provide a deeper understanding of the PA elasticity, PA hydrogels, and the ruggedness of their energy landscape, offering fundamental insights into intrinsically disordered proteins and guiding the design of novel synthetic materials with properties similar to those of tough soft materials 63, 64, 65.
@@
The structural quantities, such as absolute charge displacement and dipole formation, also highlight the mechanisms underlying hysteresis and sharp structural transitions. Our study is further comprehended by the analytical theory of the diblock PA chain, where we have employed the generalized random-phase approximation (GRPA) to estimate the system’s free energy, which also demonstrates a sharp structural transition with stretching force59. Additionally, we have extended our study of the multi-block PA chain and the biologically relevant systems, where we have investigated the elastic behavior of IDPs, such as LAF1 and DDX4. This considered system demonstrates the sharp-structural transition in the presence of the larger charged blocks’ elastic softening regimes in such systems.
@@ 2 Simulation Model
where $\varepsilon$ is the dielectric constant, which for simplicity is assumed to be homogeneous; $z_{i}=\pm e$ denotes the charge valency, and $L_{pbc}$ is the length of the primary periodic simulation box. The first summation runs over all periodic images of the primary box, except for the one that avoids self-interactions, i.e., ${\mathbf{n}}(n_{x},n_{y},n_{z})=(0,0,0)$. The long-range Coulomb interaction is computed using the Ewald summation technique 66, 67, 68, 69. The total interaction potential of the PA chain is written as $U=U_{S}+U_{LJ}+U_{C}$ and the corresponding interaction force is given by ${\bm{F}}_{i}=-\nabla_{i}U$.
@@ 3 Results
3.1 Stretching of diblock PA chain
A diblock PA chain tends to adopt a compact globular conformation even at relatively low electrostatic strengths, owing to the segregation of oppositely charged monomers into distinct blocks 56, 70, 71, 72, 62. To characterize the mechanical response of the diblock PA chain under external stretching force, we examine the root-mean-square end-to-end distance, ${R_{e}}=\sqrt{\langle({\bm{r}}_{N}-{\bm{r}}_{1})^{2}\rangle}$, as a function of the externally applied force $f$ on its terminal ends. Figure 2 shows the corresponding force–extension curve (FEC) of the chain for a range of electrostatic strengths ($\Gamma_{e}$ = 0.01–5). The solid lines represent the force-extension process, during which the chain is gradually stretched by increasing the applied force at both ends, whereas the dashed lines depict the subsequent force-relaxation process, where the applied force is progressively reduced. During the force-increasing stage, chain conformations obtained at lower forces are used as initial configurations, which are then equilibrated over a sufficiently long simulation time before analyzing $R_{e}$ and other relevant physical quantities. An identical procedure is employed during the force-relaxation process.
@@ 3.1 Stretching of diblock PA chain
3.1.1 Scaling regimes of a diblock PA chain
The force-extension curve of the PA chain can be categorized into four distinct regimes, based on distinct power-law relations between force and extension. To elucidate various scaling behaviors, we specifically analyze the force-extension response at $\Gamma_{e}=5$ as shown in Figure 3a. In the weak force limit $f/\Gamma_{e}<1$, the extension of the chain grows linearly with applied force at its ends, $R_{e}\sim f$, which is referred to as the linear response regime (LR). This is consistent with the entropic elasticity of an ideal polymer whose elasticity is given by entropic contributions 47, 73.
@@ 3.1.1 Scaling regimes of a diblock PA chain
In the limit of very high force $f/\Gamma_{e}\gg 1$, the chain nearly approaches full extension. In this regime, the extension force overpowers the entropic and electrostatic interactions. Therefore, the response follows more like a worm-like chain (WLC) scaling47, irrespective of the strength of electrostatic interactions, as shown in Figure 2 and Figure 3a. The force extension in the WLC model can be expressed
@@ 3.1.3 Elastic Softening
3.1.4 Charge Segregation and Dipole Formation
@@
3.2 Theory of Diblock PA Chain Under Extension
@@
The entropic part of free energy consists of the confinement and stretching contribution15, 74,
@@
The short-range repulsion due to the excluded volume interactions in a globule of volume $V$ is computed assuming polymer monomers are hard spheres of diameter $\sigma$ at packing fraction $\eta$. Therefore, we use the Carnahan–Starling equation of state to compute the free energy relation, which is best fitted for hard spheres at larger densities 75, 76, 77, 78. Thereby, the excluded volume contribution can be expressed as
@@
where $\eta={\pi\phi}/{6}$, $\phi={N\sigma^{3}}/{V}$, and $V=4\pi R_{g}^{3}/3$. The last term in the free energy ensures compatibility with $\Theta$-solvent conditions74.
@@
where $r_{D}=(4\pi\Gamma_{e}z_{f}\phi)^{-1/2}\sigma$ is the Debye screening length for the system of free charges and $z_{f}$ is the fraction of charged monomers. The complete details of the derivation of the free energy can be found in Ref. 74. The electrostatic correlation-free energy is calculated using the random phase approximation for the diblock PA, which requires the single-chain structure factor $g_{\mathrm{db}}(q)$ of the PA chain as input in the theory. For diblock PA chain the structure factor is approximated as ${\mathrm{g}}_{db}(q)\approx 1+{12z_{f}}/{(q\sigma)^{2}}$. Rumyantsev et. al. 74 proposed in the GRPA to exclude the non-physical modes of the charge density fluctuation by considering the cut-off wavevector $q_{0}=(9\pi^{2}z_{f}\phi)^{1/3}\sigma^{-1}$74. Integrating Eq. 19, the RPA free energy is obtained, which is also valid in the strong correlation limit,
$F_{RPA}$ $=\frac{1}{12\pi^{2}\left(\frac{r_{D}}{\sigma}\right)^{3}}\Bigg[t_{0}^{3}\ln\left(1+\frac{1}{t_{0}^{2}}+\frac{1}{s^{2}t_{0}^{4}}\right)-t_{0}-\frac{(1-\chi)^{3/2}}{\sqrt{2}}\arctan\!\left(\frac{\sqrt{2}\,t_{0}}{\sqrt{1-\chi}}\right)$ (20)
$-\frac{(1+\chi)^{3/2}}{\sqrt{2}}\arctan\!\left(\frac{\sqrt{2}\,t_{0}}{\sqrt{1+\chi}}\right)\Bigg]$
@@
3.3 Effect of charge sequence
@@
3.4 Intrinsic Disordered Proteins
To further elucidate the role of charge sequencing, we examine two intrinsically disordered proteins (IDPs), LAF1 and DDX4. LAF1 is localized in P-granules and plays a crucial role in liquid–liquid phase separation and phytochrome-mediated signaling pathways in plant cells, whereas DDX4 is implicated in antiviral defense mechanisms and translation regulation. We employ coarse-grained representations of these proteins comprising $N=168$ and $N=236$ residues for LAF1 and DDX4, respectively. Detailed descriptions of the simulation methodology and sequence information are provided in the Supplementary Material 79.
@@ 4 Conclusions
Our study reveals hysteresis in the force–extension and relaxation process indicating enhanced energy dissipation —a hallmark of viscoelastic materials such as hydrogels and biopolymer networks 80, 81, 82, 83, 84, 85, 86. In such systems, hysteresis typically arises from association and dissociation of physical bonds and the redistribution of mechanical load, contributing to energy loss during cyclic deformation. The force-induced softening has been observed in biomolecular systems such as dsDNA and proteins, where overstretching destabilizes secondary structural organization, especially hydrogen bonding, thereby inducing significant energy dissipation and mechanical softening80, 81, 82, 83, 84, 85, 86.
Analysis of the force–extension curves (FECs) reveals four distinct elastic regimes: an initial linear response followed by three nonlinear regions characterized by the elastic modulus. In the low-force regime, the modulus remains constant $E\sim f^{0}$. At intermediate forces, the chain exhibits stress stiffening $E\sim f^{1.2}$, which is subsequently followed by a stress-softening regime $E\sim\exp(-\alpha_{0}f/\Gamma_{e})$, arising from strong electrostatic coupling. This stress-softening regime is absent at low electrostatic strengths but becomes prominent as block length increases. At higher forces, the chain stiffens again with $E\sim f^{3/2}$ 47, 52. The stress-softening behavior is predominantly associated with block PA chains, whereas such effects are generally not observed in neutral polymers or uniformly charged polyelectrolytes. Our analysis reveals that the microscopic origin of the exponential softening is due to electrostatic coupling between oppositely charged folded domains of the double hairpin-like structures. Using structural inputs from this regime, we analytically show that the elastic modulus decays exponentially with applied force. The theoretical framework aligns closely with the simulation behaviors, capturing both continuous transitions in the weakly correlated regime and first-order-like discontinuous transitions under strong electrostatic correlations, thereby validating the simulation findings.
@@
In summary, our study reveals new aspects of polymer mechanics resulting from charge segregation and electrostatic correlations. The mechanical response of a single polymer has paramount relevance in understanding the behavior of polymeric materials, brushes, networks15, 87, protein functionality, RNA energy landscapes16, and cellular elasticity88, 18. This provides a direct link between macroscopic elastic behavior and the microscopic structural arrangements of the chain18. More broadly, our findings provide implications, ranging from the mechanical characterization of biological tissues to the rational design of responsive self-healing hydrogels89, 90, 65, tunable soft materials of low-dissipation91, 92, etc.
{acknowledgement}
RP and SPS acknowledge financial support from the UGC-India and DST-SERB Grant No. CRG/2020/000661, and computational time at IISER Bhopal and the Param Himalaya NSM facility.

### v2

# Force-induced Elastic Softening and Conformational Transitions in a Polyampholyte chain

@@ Abstract
The mechanical response of intrinsically disordered proteins (IDPs) and polyampholyte (PA) chains is vital for understanding their biological functions and designing functional materials. We investigate the force-extension behavior of a PA chain with distinct charge sequences using molecular dynamics simulations and a theoretical approach based on the generalized random-phase approximation (GRPA). A diblock PA chain under extensional force undergoes a continuous coil-to-stretch transition at weak electrostatic coupling, which sharpens into a globule-coil-like transition at stronger coupling. The GRPA theory quantitatively captures these behaviors, including the sharp conformational transition and its dependence on electrostatic strength. Simulations reveal pronounced hysteresis during the force-extension and relaxation processes. Additionally, the elastic modulus exhibits four regimes: an initial plateau, stress stiffening, an exponential stress-softening behavior, and a stress stiffening regime. Using the theoretical model and structural input of the PA chain, we have demonstrated that the elastic modulus in the elastic softening regime decreases exponentially, $E\sim\exp(-\alpha_{0}f/\Gamma_{e})$, as a function of $f$, which aligns with the simulation results. The elastic response of the PA chain is further examined across different charge sequences, where both elastic softening and sharp transitions are absent at smaller block lengths. Finally, coarse-grained models of IDPs such as LAF-1 and DDX4 exhibit similar nonlinear elasticity, highlighting the universality of these mechanisms. Our results establish a fundamental link between electrostatic correlations, charge sequence, and nonlinear elasticity, bridging molecular interactions and macroscopic mechanics.
I Introduction
@@
A wide variety of biopolymers—including nucleic acids, and intrinsically disordered proteins (IDPs)—as well as synthetic macromolecular systems, contain both acidic and basic functional groups [1, 2, 3]. These systems play essential roles in numerous biological processes, such as cell signaling [4, 5], transcriptional regulation [6, 7], muscle elasticity [8], liquid–liquid phase separation (LLPS), and the formation of biomolecular condensates [9, 10, 11, 12, 13, 14]. Understanding their mechanical response is crucial for elucidating the behavior of biological networks [15], protein functionality [4, 5, 6, 7], and the folding energy landscapes of RNA and DNA [16, 17, 18, 19]. Many such polymers and proteins possess folded structural domains that undergo cooperative unfolding under applied forces, resulting in complex force–extension curves (FECs) characterized by abrupt extensions and nonlinear mechanical responses [20, 21, 22, 23, 24].
@@
The theoretical frameworks for the elastic response of a polymer have led to interesting scaling relationships that describe mechanical behavior across different regimes[25, 15, 26, 27, 28, 25, 29, 18, 30]. These relationships have been supported by both computational approaches and experimental observations[18, 31, 32, 33, 34, 35, 36, 20, 21, 22, 23, 24]. The single‑molecule techniques such as atomic force microscopy (AFM) and optical/magnetic tweezers have significantly transformed the study of the polymer’s elastic behavior[37, 38, 17, 34, 16, 39, 40, 41, 42, 18, 34, 43]. These methods enable direct measurement of force–extension relationships at the level of individual chains, thereby validating theoretical predictions, identifying distinct mechanical regimes, and highlighting how molecular heterogeneity influences bulk mechanical behavior.
@@
The elastic response of a linear, uncharged polymer in the low-force regime is characterized by purely entropic behavior, with the force–extension curve being linear, and the elastic modulus ($E$) remaining constant. At large stretching forces, a nonlinear relationship emerges between force and extension. In the case of the freely jointed chain (FJC) model[44], the extension approaches $R_{e}>\approx R_{max}[1-k_{B}T/f\ell_{K}]$, where $R_{e}$, $R_{max}$, $k_{B}T$, and $\ell_{K}$ are the chains extension, contour length, thermal energy, and Kuhn length, respectively. Alternatively, for semiflexible chains such as double-stranded DNA, a worm-like chain (WLC) model developed by Marko and Siggia provides the force extension as $R_{e}\approx R_{max}[1-\sqrt{k_{B}T/2f\ell_{K}]}$ [25, 45, 38, 46, 19, 47, 48]. The WLC model provides a fairly good description of the elastic response of ssRNA, ssDNA, PolyEthylene Glycol, and various other proteins across a broad range of forces[34, 49, 50, 18, 51, 15]. In these large-force limits, the elastic modulus scales as $E\sim f^{2}$ for the FJC and $E\sim f^{1.5}$ for the WLC model[52]. More importantly, in the intermediate regimes, the force extension typically depends on the nature of the interactions among various domains of the polymer, excluded volume, electrostatic interactions, and ionic concentrations. These interactions can lead to distinct elastic behaviors.
@@
Previous investigations have examined the force–extension behavior of neutral polymers, polyelectrolytes, and zwitterionic polymers [35, 53, 54, 32, 33, 55]. These models are typically formulated for homogeneously charged chains and therefore do not capture the complex electrostatic environments that can give rise to the rugged energy landscapes often observed in proteins. Despite extensive work on neutral and homogeneously charged polymers, the role of electrostatic sequence heterogeneity and charge segregation in determining the nonlinear elastic response, hysteresis, and softening behavior of polyampholytes (PAs) remains poorly understood. PA serves as a simplified model system for intrinsically disordered proteins (IDPs) and other charged proteins. Their conformational and phase behaviors depend intricately on parameters such as net charge, charge sequence or block length, chain rigidity, and ionic concentration [56, 57, 58, 59, 60, 61, 62]. The influence of charge sequence and the strength of electrostatic interactions on their elastic behavior remains crucial yet not explored. Addressing this gap could provide a deeper understanding of the PA elasticity, PA hydrogels, and the ruggedness of their energy landscape, offering fundamental insights into intrinsically disordered proteins and guiding the design of novel synthetic materials with properties similar to those of tough soft materials [63, 64, 65].
@@
The structural quantities, such as absolute charge displacement and dipole formation, also highlight the mechanisms underlying hysteresis and sharp structural transitions. Our study is further comprehended by the analytical theory of the diblock PA chain, where we have employed the generalized random-phase approximation (GRPA) to estimate the system’s free energy, which also demonstrates a sharp structural transition with stretching force[59]. Additionally, we have extended our study of the multi-block PA chain and the biologically relevant systems, where we have investigated the elastic behavior of IDPs, such as LAF1 and DDX4. This considered system demonstrates the sharp-structural transition in the presence of the larger charged blocks’ elastic softening regimes in such systems.
@@ II Simulation Model
where $\varepsilon$ is the dielectric constant, which for simplicity is assumed to be homogeneous; $z_{i}=\pm e$ denotes the charge valency, and $L_{pbc}$ is the length of the primary periodic simulation box. The first summation runs over all periodic images of the primary box, except for the one that avoids self-interactions, i.e., ${\mathbf{n}}(n_{x},n_{y},n_{z})=(0,0,0)$. The long-range Coulomb interaction is computed using the Ewald summation technique [66, 67, 68, 69]. The total interaction potential of the PA chain is written as $U=U_{S}+U_{LJ}+U_{C}$ and the corresponding interaction force is given by ${\bm{F}}_{i}=-\nabla_{i}U$.
@@ III Results
III.1 Stretching of diblock PA chain
A diblock PA chain tends to adopt a compact globular conformation even at relatively low electrostatic strengths, owing to the segregation of oppositely charged monomers into distinct blocks [56, 70, 71, 72, 62]. To characterize the mechanical response of the diblock PA chain under external stretching force, we examine the root-mean-square end-to-end distance, ${R_{e}}=\sqrt{\langle({\bm{r}}_{N}-{\bm{r}}_{1})^{2}\rangle}$, as a function of the externally applied force $f$ on its terminal ends. Figure 2 shows the corresponding force–extension curve (FEC) of the chain for a range of electrostatic strengths ($\Gamma_{e}$ = 0.01–5). The solid lines represent the force-extension process, during which the chain is gradually stretched by increasing the applied force at both ends, whereas the dashed lines depict the subsequent force-relaxation process, where the applied force is progressively reduced. During the force-increasing stage, chain conformations obtained at lower forces are used as initial configurations, which are then equilibrated over a sufficiently long simulation time before analyzing $R_{e}$ and other relevant physical quantities. An identical procedure is employed during the force-relaxation process.
@@ III.1 Stretching of diblock PA chain
III.1.1 Scaling regimes of a diblock PA chain
The force-extension curve of the PA chain can be categorized into four distinct regimes, based on distinct power-law relations between force and extension. To elucidate various scaling behaviors, we specifically analyze the force-extension response at $\Gamma_{e}=5$ as shown in Figure 3a. In the weak force limit $f/\Gamma_{e}<1$, the extension of the chain grows linearly with applied force at its ends, $R_{e}\sim f$, which is referred to as the linear response regime (LR). This is consistent with the entropic elasticity of an ideal polymer whose elasticity is given by entropic contributions [47, 73].
@@ III.1.1 Scaling regimes of a diblock PA chain
In the limit of very high force $f/\Gamma_{e}\gg 1$, the chain nearly approaches full extension. In this regime, the extension force overpowers the entropic and electrostatic interactions. Therefore, the response follows more like a worm-like chain (WLC) scaling[47], irrespective of the strength of electrostatic interactions, as shown in Figure 2 and Figure 3a. The force extension in the WLC model can be expressed
@@ III.1.3 Elastic Softening
III.1.4 Charge Segregation and Dipole Formation
@@
III.2 Theory of Diblock PA Chain Under Extension
@@
The entropic part of free energy consists of the confinement and stretching contribution[15, 74],
@@
The short-range repulsion due to the excluded volume interactions in a globule of volume $V$ is computed assuming polymer monomers are hard spheres of diameter $\sigma$ at packing fraction $\eta$. Therefore, we use the Carnahan–Starling equation of state to compute the free energy relation, which is best fitted for hard spheres at larger densities [75, 76, 77, 78]. Thereby, the excluded volume contribution can be expressed as
@@
where $\eta={\pi\phi}/{6}$, $\phi={N\sigma^{3}}/{V}$, and $V=4\pi R_{g}^{3}/3$. The last term in the free energy ensures compatibility with $\Theta$-solvent conditions[74].
@@
where $r_{D}=(4\pi\Gamma_{e}z_{f}\phi)^{-1/2}\sigma$ is the Debye screening length for the system of free charges and $z_{f}$ is the fraction of charged monomers. The complete details of the derivation of the free energy can be found in Ref. [74]. The electrostatic correlation-free energy is calculated using the random phase approximation for the diblock PA, which requires the single-chain structure factor $g_{\mathrm{db}}(q)$ of the PA chain as input in the theory. For diblock PA chain the structure factor is approximated as ${\mathrm{g}}_{db}(q)\approx 1+{12z_{f}}/{(q\sigma)^{2}}$. Rumyantsev et. al. [74] proposed in the GRPA to exclude the non-physical modes of the charge density fluctuation by considering the cut-off wavevector $q_{0}=(9\pi^{2}z_{f}\phi)^{1/3}\sigma^{-1}$[74]. Integrating Eq. 19, the RPA free energy is obtained, which is also valid in the strong correlation limit,
$F_{RPA}=\frac{1}{12\pi^{2}\left(\frac{r_{D}}{\sigma}\right)^{3}}\Bigg[t_{0}^{3}\ln\left(1+\frac{1}{t_{0}^{2}}+\frac{1}{s^{2}t_{0}^{4}}\right)-t_{0}-\frac{(1-\chi)^{3/2}}{\sqrt{2}}\arctan\!\left(\frac{\sqrt{2}\,t_{0}}{\sqrt{1-\chi}}\right)\\ -\frac{(1+\chi)^{3/2}}{\sqrt{2}}\arctan\!\left(\frac{\sqrt{2}\,t_{0}}{\sqrt{1+\chi}}\right)\Bigg]$ (20)
@@
III.3 Effect of charge sequence
@@
III.4 Intrinsic Disordered Proteins
To further elucidate the role of charge sequencing, we examine two intrinsically disordered proteins (IDPs), LAF1 and DDX4. LAF1 is localized in P-granules and plays a crucial role in liquid–liquid phase separation and phytochrome-mediated signaling pathways in plant cells, whereas DDX4 is implicated in antiviral defense mechanisms and translation regulation. We employ coarse-grained representations of these proteins comprising $N=168$ and $N=236$ residues for LAF1 and DDX4, respectively. Detailed descriptions of the simulation methodology and sequence information are provided in the Supplementary Material [79].
@@ IV Conclusions
Our study reveals hysteresis in the force–extension and relaxation process indicating enhanced energy dissipation —a hallmark of viscoelastic materials such as hydrogels and biopolymer networks [80, 81, 82, 83, 84, 85, 86]. In such systems, hysteresis typically arises from association and dissociation of physical bonds and the redistribution of mechanical load, contributing to energy loss during cyclic deformation. The force-induced softening has been observed in biomolecular systems such as dsDNA and proteins, where overstretching destabilizes secondary structural organization, especially hydrogen bonding, thereby inducing significant energy dissipation and mechanical softening[80, 81, 82, 83, 84, 85, 86].
Analysis of the force–extension curves (FECs) reveals four distinct elastic regimes: an initial linear response followed by three nonlinear regions characterized by the elastic modulus. In the low-force regime, the modulus remains constant $E\sim f^{0}$. At intermediate forces, the chain exhibits stress stiffening $E\sim f^{1.2}$, which is subsequently followed by a stress-softening regime $E\sim\exp(-\alpha_{0}f/\Gamma_{e})$, arising from strong electrostatic coupling. This stress-softening regime is absent at low electrostatic strengths but becomes prominent as block length increases. At higher forces, the chain stiffens again with $E\sim f^{3/2}$ [47, 52]. The stress-softening behavior is predominantly associated with block PA chains, whereas such effects are generally not observed in neutral polymers or uniformly charged polyelectrolytes. Our analysis reveals that the microscopic origin of the exponential softening is due to electrostatic coupling between oppositely charged folded domains of the double hairpin-like structures. Using structural inputs from this regime, we analytically show that the elastic modulus decays exponentially with applied force. The theoretical framework aligns closely with the simulation behaviors, capturing both continuous transitions in the weakly correlated regime and first-order-like discontinuous transitions under strong electrostatic correlations, thereby validating the simulation findings.
@@
In summary, our study reveals new aspects of polymer mechanics resulting from charge segregation and electrostatic correlations. The mechanical response of a single polymer has paramount relevance in understanding the behavior of polymeric materials, brushes, networks[15, 87], protein functionality, RNA energy landscapes[16], and cellular elasticity[88, 18]. This provides a direct link between macroscopic elastic behavior and the microscopic structural arrangements of the chain[18]. More broadly, our findings provide implications, ranging from the mechanical characterization of biological tissues to the rational design of responsive self-healing hydrogels[89, 90, 65], tunable soft materials of low-dissipation[91, 92], etc.
Acknowledgments
RP and SPS acknowledge financial support from the UGC-India and DST-SERB Grant No. CRG/2020/000661, and computational time at IISER Bhopal and the PARAM Smriti NSM facility.